OpenStudy (shes.radical):

Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). A.) Y = 1over3x + 7 B.) Y = -3x + 3 C.) Y = 1over3x + 3 D.) Y = -3x + 7

OpenStudy (campbell_st):

what is the slope of the given line...?

OpenStudy (shes.radical):

No Clue . Thats the entire question .

OpenStudy (campbell_st):

well its in the information. The slope intercept form of a line is \[y = mx + b\] m is the slope and b is the y intercept. so looking at th line you have been given, what is the slope..?

OpenStudy (shes.radical):

-3

OpenStudy (campbell_st):

great... so you can use that information the line is y = -3x + b so you can eliminate 2 choices

OpenStudy (campbell_st):

next use the point (-1, 6) and y = -3x + b this will help you find b, so substitute x = -1 and y = 6 to find b \[6 = - 3 \times -1 + b\] what's the value of b..?

OpenStudy (shes.radical):

3...?

OpenStudy (campbell_st):

great the line you need is y = -3x + 3 is that in your list...?

OpenStudy (shes.radical):

yes (: thanks you

OpenStudy (shes.radical):

and ! @campbell_st i have other questions similar to this one but i can do them now (: but some other ones say "perpendicular" instead of "parallel" .. is there sumthing that i must do different ?

OpenStudy (campbell_st):

well if you post them as new questions there a lots of people who can help... but remember y = mx + b m = slope.... substitute the point to find b and the solution.

OpenStudy (shes.radical):

alrighty, thanks